# iSEARLE: Improving Textual Inversion for Zero-Shot Composed Image Retrieval

### v1

@@ III-B Textual Inversion Network $\bm{\phi}$ Pre-training
OTI proves effective in generating pseudo-words that not only capture the visual information of an image but also interact fruitfully with actual words. However, its iterative and optimization-based nature results in a non-negligible amount of time for its execution. To address this issue, we propose an approach for training a textual inversion network $\phi$ capable of predicting the pseudo-word tokens in a single forward pass by distilling knowledge from a collection of OTI pre-generated tokens. In other words, $\phi$ serves as a more efficient surrogate model of OTI, offering a faster and computationally less demanding approximation. The right part of Fig. 2 illustrates an overview of the pre-training phase.
We aim to obtain a single model capable of inverting images from any domain, without the requirement of labeled training data. Specifically, we design an MLP-based textual inversion network $\phi$ with three linear layers, each followed by a GELU [55] activation function and a dropout layer.
Starting from an unlabeled pre-training dataset $\mathcal{D}$, we apply OTI to each image. Despite being time-intensive, this step is only required once, which makes the process acceptable. This results in a collection of pseudo-word tokens denoted as $\overline{\mathcal{V}}_{*}=\{{\bar{v}_{*}^{j}}\}_{j=1}^{N}$, with $N$ representing the total number of images in $\mathcal{D}$. Our goal is to distill the knowledge captured by OTI in $\overline{\mathcal{V}}_{*}$ to $\phi$. Given an image $I\in\mathcal{D}$, we extract its features via the CLIP visual encoder, resulting in $x=\psi_{I}(I)$. We exploit $\phi$ to predict the pseudo-word token $v_{*}=\phi(x)$. We minimize the distance between the predicted pseudo-word token $v_{*}$ and the associated pre-generated token $\bar{v}_{*}\in\overline{\mathcal{V}}_{*}$ while maximizing the discriminability of each token. To achieve this, we rely on a symmetric contrastive loss inspired by SimCLR [56, 42] as follows:
@@ IV-A Triplets Annotation
CIRCO is based on images sourced from COCO 2017 [17] unlabeled set, which encompasses 123,403 images. This dataset was suitable for our goals as it comprises open-domain real-world images that portray a wide range of subjects. In addition, we opted for the COCO unlabeled set over the training one to avoid any pre-existing model biases, as the latter is commonly used for pre-training. In the COCO labeled sets, each object in an image is categorized under one of 12 supercategories: person, animal, sports, vehicle, food, accessory, electronic, kitchen, furniture, indoor, outdoor, and appliance.
@@
The annotation tool selects a reference image at random and presents it alongside a gallery of 50 candidate target images. The target images must be visually similar to the reference one yet exhibit discernible disparities, as CIR requires the differences between them to be describable with a relative caption. Consequently, we choose the candidate target images based on their visual similarity to the reference image as per the CLIP features. To prevent the inclusion of near-identical images, we exclude those with a cosine similarity exceeding 0.92. The annotators are allowed to skip the current reference image if no suitable target is found in the gallery. On the contrary, when a suitable target image is available, the annotator selects it and writes the shared concept, which represents the common characteristics between the reference and target images. We collect the shared concept to address any potential ambiguities. Finally, the annotator crafts a relative caption from the prefix “Unlike the provided image, I want a photo of {shared concept} that”. Since our goal is to create a challenging dataset comprising truly relative captions, we ensure that they are formulated in such a way that avoids references to subjects mentioned in the shared concept. In this way, the subject of the relative caption needs to be deduced from the reference image alone.
At the end of this phase, we obtain 1020 triplets composed of a reference image, a relative caption, and a single target image.
@@ IV-B Multiple Ground Truths Annotation
We propose to facilitate the annotation process by exploiting our approach to retrieve the images from which the ground truths are selected. In particular, we employ SEARLE to generate the pseudo-word $S_{*}$ corresponding to the reference image. Then, we carry out text-to-image retrieval based on the query ‘a photo of {shared concept} $S_{*}$ that {relative caption}”. During the annotation phase, we incorporate the shared concept into the query because it improves performance. Indeed, considering the single ground truth triplets obtained in Sec. IV-A, we achieve a Recall@$100$ of 82.15 with the shared concept and of 66.25 without it. In the gallery of images used for selecting the multiple ground truths, the annotation tool presents the top 100 retrieved images using our approach, along with the top 50 images most visually similar to the target one.
At the end of this phase, we have 4624 ground truths, of which 4097 were retrieved employing our method and 527 using the similarity with the target image. Since SEARLE achieves a Recall@$100$ of 82.15, by approximation, we estimate that about 82.15% of the total ground truths are present in the top 100 retrieved images. Consequently, the estimated total number of ground truths in our dataset is approximately $4097/0.8215\approx 4,987$. Given that we labeled 4624 images as ground truth, we can infer that the annotated ones are $4,624/4,987\approx 92.7\%$ of the total. Therefore, we estimate that the proposed annotation strategy allows us to reduce the percentage of missing ground truths in the dataset to less than 10%.
Thanks to this second annotation step, we labeled additional $4624\!-\!1020\!=\!3604$ ground truths that would have otherwise been considered false negatives. Furthermore, this phase enables us to estimate the percentage of missing ground truths within the dataset. Notably, this estimation is unfeasible for CIR datasets featuring only a single ground truth, such as FashionIQ [14] and CIRR [2], as they lack any information about the total number of ground truths. Indeed, in these datasets, the annotation process concludes upon the completion of the triplet construction.
@@ IV-C Semantic Aspects Annotation
To allow a fine-grained semantic analysis on CIRCO, we perform a third annotation phase to assign semantic aspects to each query depending solely on its relative caption. In particular, we consider the same semantic categories as CIRR [2], such as “direct addressing” and “compare & change”.
@@ IV-D Dataset Analysis
CIRCO comprises 1020 queries, randomly divided into 220 and 800 for the validation and test set, respectively. The total number of ground truths is 4624, 4.53 per query on average. Figure 5 shows a histogram representing the number of queries per number of ground truths. The maximum number of ground truths annotated for a query is 21, while the modal value is 2.
Figure 5: Histogram of the number of queries per number of ground truths for the CIRCO dataset.
@@ IV-E Evaluation Metric
Thanks to the reduced false negatives and multiple ground truths of CIRCO, we can rely on a more fine-grained metric for performance evaluation, such as mean Average Precision (mAP). Indeed, mAP considers also the ranks in which the ground truths are retrieved. Specifically, we compute mAP@$K$, with K ranging from 5 to 50, as follows:
@@
where $N$ is the number of queries, $G_{n}$ is the number of ground truths of the $n$-th query, $P@k$ is the precision at rank $k$, rel@$k$ is a relevance function. The relevance function is an indicator function that equals 1 if the image at rank $k$ is labeled as a ground truth and equals 0 otherwise.
@@ V Experimental Results
We present two variants of our method: iSEARLE, based on CLIP ViT-B/32, and iSEARLE-XL, using CLIP ViT-L/14 as the backbone. From now on, we refer to ViT-B/32 and ViT-L/14 as B/32 and L/14, respectively.
@@ V-A Implementation Details
Regarding the Optimization-based Textual Inversion (OTI), we perform 500 iterations with a learning rate of $22e\!-\!$ 2. We set the loss weights $\lambda_{content}$ and $\lambda_{OTIgpt}$ in Eq. 3 to 1 and 0.5, respectively. We set the standard deviation $\gamma$ of the Gaussian noise to 0.64 and 0.16 respectively for iSEARLE-OTI and iSEARLE-XL-OTI. For the textual inversion network $\phi$, we train for 115 epochs, with a learning rate of $41e\!-\!$ 4 and a batch size of 256. We set the loss weights $\lambda_{distil}$ and $\lambda_{\phi gpt}$ in Eq. 6 to 1 and 0.75, respectively. The loss weight $\lambda_{pen}$ is equal to $33e\!-\!$ 3 and $21e\!-\!$ 2 for the B/32 and L/14 backbones, respectively. The temperature $\tau$ in Sec. III-B is set to 0.25. The parameter $\alpha$ of the hard negative sampling strategy (Sec. III-B) is set to 0.5. For both OTI and $\phi$, we employ the AdamW optimizer [64] with weight decay 0.01. We use an exponential moving average of 0.99 and 0.999 decay for OTI and $\phi$, respectively. During OTI we set the number of concept words $k$ associated with each image to 15, while during the training of $\phi$ to 150. We tune each hyperparameter individually with a grid search on the CIRR validation set.
Using a single NVIDIA A100 GPU, OTI for iSEARLE-XL takes $\sim$35 seconds for a single image and $\sim$1.1 seconds per image with batch size 256. The training of $\phi$ for iSEARLE-XL takes 12 hours in total on a single A100 GPU. Throughout all the experiments, we adopt the pre-processing technique introduced in [5]. For retrieval, we normalize both the query and index set features to have a unit $L_{2}$-norm.
To generate the phrases used for the regularization with $\mathcal{L}_{gpt}$, we exploit the GPT-Neo-2.7B model with 2.7 billion parameters developed by EleutherAI. For each of the 20,932 class names of the Open Images V7 dataset [54], we generate 256 phrases a priori with a temperature of 0.5, with a maximum length constraint of 35 tokens. The whole process requires about 12 hours to execute on a single A100 GPU. Since this operation only needs to be performed once, the time requirements are manageable.
We employ the same set of templates at training and inference time. Since the FashionIQ dataset provides two relative captions for each triplet, at inference time, we concatenate them using the conjunction “and”. To ensure our method remains independent of the order of the concatenation, we use both possible concatenation orders and average the resulting features.
@@ V-B Quantitative Results
We provide the results of both iSEARLE and iSEARLE-OTI. We compare our method with several zero-shot baselines: 1) Text-only: we compute the similarity using only the CLIP features of the relative caption; 2) Image-only: we retrieve the most similar images to the reference one; 3) Image + Text: we sum together the CLIP features of the reference image and the relative caption; 4) Captioning: we substitute the pseudo-word token with the caption of the reference image obtained via a pre-trained captioning model [65] [^4] https://huggingface.co/laion/CoCa-ViT-B-32-laion2B-s13B-b90k. In addition, we compare the proposed approach with the conference version of this work [9] and state-of-the-art ZS-CIR methods: Pic2Word [18], Context-I2W [36], and LinCIR [37]. For a fair comparison, we only consider competing methods that rely on the CLIP model without fine-tuning its weights and that, at inference time, do not require any additional pre-trained models, such as LLMs.
TABLE III: Quantitative results on CIRR test set. Best and second-best scores are highlighted in bold and underlined, respectively. indicates results from the original paper. – denotes results not reported in the original paper.
Recall@$K$
@@
TABLE IV: Quantitative results on CIRCO test set. Best and second-best scores are highlighted in bold and underlined, respectively. indicates results from the original paper.
mAP@$K$
@@
TABLE V: Quantitative results on CIRCO test set for each semantic category. Best and second-best scores are highlighted in bold and underlined, respectively.
@@
CIRCO In Table IV we report the results for the CIRCO test set. Firstly, we observe that, in contrast to FashionIQ and CIRR, Image+Text outperforms Image-only and Text-only. This result indicates that CIRCO contains queries where both the reference image and the relative caption are equally crucial for retrieving the target images. Secondly, iSEARLE achieves a considerable improvement over all the baselines and even outperforms Pic2Word, despite a smaller backbone. Considering the L/14 backbone, iSEARLE-XL would achieve the best results. However, we recall that the conference version of our method SEARLE-XL [9] was used to ease the annotation process of CIRCO. Therefore, since the core of the approach proposed in this work is the same, it is likely that the results could exhibit some sort of bias. Still, we report them for completeness.
@@
Domain Conversion Table VI illustrates the results for the domain conversion task. Following [18], the query images are sourced from 200 classes of ImageNet [61] validation set while the target ones belong to ImageNet-R [63]. We recall that we relied on the ImageNet unlabeled test set as the pre-training dataset of our method, so there is no overlap with the evaluation images of the domain conversion task. The purpose of this experiment is to study how our model can convert the domain of a query image by using the prompt “$\{domain\}$ of $S_{*}$”, where $\{domain\}$ is a word that indicates the domain, e.g. toy or origami. We consider the retrieved image correct if its class is the same as that of the query image and its domain matches the one specified by the prompt. For instance, given the domain “toy” and a query image containing a real-world shark, the goal is to retrieve images depicting a shark toy. The results show that, both for the B/32 and L/14 backbones, our method outperforms all the baselines. This outcome highlights that the proposed approach has better generalization and adaptability capabilities to tasks different from standard composed image retrieval.
TABLE VI: Quantitative results for the domain conversion task. The query images are from the ImageNet validation set while the target ones belong to ImageNet-R. Best and second-best scores are highlighted in bold and underlined, respectively. indicates results from the original paper.
@@
Object Composition We provide the results for the object composition task on the COCO validation set [17] in Tab. VII. This task aims to retrieve an image comprising an object specified with a single query image and other objects described with text. The object composition task closely resembles the personalized retrieval one [42], differing mainly in that the latter involves queries composed of multiple images depicting the same object instance. Following [18], we use the prompt “a photo of $S_{*}$, $\{obj_{1}\}$ and $\{obj_{2}\}$, …, and $\{obj_{n}\}$”, where $\{obj_{i}\}$ are text descriptions of objects, e.g. mouse, laptop or kite. For both the B/32 and L/14 backbones, we notice that our approach achieves state-of-the-art results. Considering the L/14 backbone, we observe a performance gap between the OTI and $\phi$ variants of the proposed method. We suppose this result is due to the similarity between the object composition and personalized retrieval tasks, where it has been shown that an optimization-based method is more suitable and achieves better performance [42]. The results obtained by PALAVRA [42] on the object composition task confirm our hypothesis, as it achieves significantly better relative performance than in composed image retrieval, e.g. on CIRCO. In addition, we observe that LinCIR [37] obtains considerably worse performance than iSEARLE-XL. We suppose this outcome is due to their language-only training strategy, which makes their model struggle to capture fine-grained visual details.
TABLE VII: Quantitative results for the object composition task on COCO validation set. Best and second-best scores are highlighted in bold and underlined, respectively. indicates results from the original paper.
Backbone Method R@$1$ R@$5$ R@$10$

### v2

@@ III-B Textual Inversion Network $\bm{\phi}$ Pre-training
OTI proves to be effective in generating pseudo-words that not only capture the visual information of an image but also interact fruitfully with actual words. However, its iterative and optimization-based nature results in a non-negligible amount of time for its execution (see Sec. V-A). To address this issue, we propose an approach for training a textual inversion network $\phi$ capable of predicting the pseudo-word tokens in a single forward pass by distilling knowledge from a collection of OTI pre-generated tokens. In other words, $\phi$ serves as a more efficient surrogate model of OTI, offering a faster and computationally less demanding approximation. The right part of Fig. 2 illustrates an overview of the pre-training phase.
We aim to obtain a single model capable of inverting images from any domain without the requirement of labeled training data. Specifically, we design an MLP-based textual inversion network $\phi$ with three linear layers, each followed by a GELU [52] activation function and a dropout layer.
Starting from an unlabeled pre-training dataset $\mathcal{D}$, we apply OTI to each image. Although this step is time-intensive, it is only required once, making it acceptable. This results in a collection of pseudo-word tokens, denoted as $\overline{\mathcal{V}}_{*}=\{{\bar{v}_{*}^{j}}\}_{j=1}^{N}$, where $N$ is the total number of images in $\mathcal{D}$. Our goal is to distill the knowledge captured by OTI in $\overline{\mathcal{V}}_{*}$ to $\phi$. Given an image $I\in\mathcal{D}$, we extract its features via the CLIP visual encoder, resulting in $x=\psi_{I}(I)$. We exploit $\phi$ to predict the pseudo-word token $v_{*}=\phi(x)$. We minimize the distance between the predicted pseudo-word token $v_{*}$ and the associated pre-generated token $\bar{v}_{*}\in\overline{\mathcal{V}}_{*}$ while maximizing the discriminability of each token. To achieve this, we employ a symmetric contrastive loss inspired by SimCLR [53, 38]:
@@ IV-A Triplets Annotation
CIRCO is based on images sourced from COCO 2017 [17] unlabeled set, which comprises 123,403 images. This dataset was suitable for our goals as it comprises open-domain real-world images that portray a wide range of subjects. In addition, we opted for the COCO unlabeled set over the training one to avoid any pre-existing model biases, as the latter is commonly used for pre-training. In the COCO labeled sets, each object in an image is categorized under one of 12 supercategories: person, animal, sports, vehicle, food, accessory, electronic, kitchen, furniture, indoor, outdoor, and appliance.
@@
The annotation tool selects a reference image at random and presents it alongside a gallery of 50 candidate target images. The target images must be visually similar to the reference one yet exhibit discernible disparities, as CIR requires the differences between them to be describable with a relative caption. Consequently, we choose the candidate target images based on their visual similarity to the reference image as per the CLIP features. To prevent the inclusion of near-identical images, we exclude those with a cosine similarity exceeding 0.92. The annotators are allowed to skip the current reference image if no suitable target is found in the gallery. On the contrary, when a suitable target image is available, the annotator selects it and writes the shared concept, which represents the common characteristics between the reference and target images. We collect the shared concept to address any potential ambiguities. Finally, the annotator crafts a relative caption from the prefix ``Unlike the provided image, I want a photo of {shared concept} that". Since our goal is to create a challenging dataset comprising truly relative captions, we ensure that they are formulated in such a way that avoids references to subjects mentioned in the shared concept. In this way, the subject of the relative caption needs to be deduced from the reference image alone.
At the end of this phase, we obtain 1020 triplets comprising a reference image, a relative caption, and a single target image.
@@ IV-B Multiple Ground Truths Annotation
We propose to facilitate the annotation process by exploiting our approach to retrieve the images from which the ground truths are selected. In particular, we employ SEARLE to generate the pseudo-word $S_{*}$ corresponding to the reference image. Then, we carry out text-to-image retrieval based on the query `a photo of {shared concept} $S_{*}$ that {relative caption}". During the annotation phase, we incorporate the shared concept into the query because it improves performance. Indeed, considering the single ground truth triplets obtained in Sec. IV-A, we achieve a Recall@100@100@ 100 of 82.15 with the shared concept and of 66.25 without it. In the gallery of images used for selecting the multiple ground truths, the annotation tool presents the top 100 retrieved images using our approach, along with the top 50 images most visually similar to the target one.
[…]
Thanks to this second annotation step, we labeled additional 46241020=36044624\!-\!1020\!=\!36044624 - 1020 = 3604 ground truths that would have otherwise been considered false negatives. Furthermore, this phase enables us to estimate the percentage of missing ground truths within the dataset. Notably, this estimation is unfeasible for CIR datasets featuring only a single ground truth, such as FashionIQ [14] and CIRR [2], as they lack any information about the total number of ground truths. Indeed, in these datasets, the annotation process concludes upon the completion of the triplet construction.
@@ IV-C Semantic Aspects Annotation
To allow a fine-grained semantic analysis on CIRCO, we perform a third annotation phase to assign semantic aspects to each query depending solely on its relative caption. In particular, we consider the same semantic categories as CIRR [2], such as ``direct addressing" and ``compare & change".
@@ IV-D Dataset Analysis
CIRCO comprises 1020 queries, randomly divided into 220 and 800 for the validation and test set, respectively. The total number of ground truths is 4624, i.e. 4.53 per query on average. The maximum number of ground truths annotated for a query is 21, while the modal value is 2.
@@ IV-E Evaluation Metric
[…]
@@
[…]
@@ V Experimental Results
We present two variants of our method: iSEARLE, based on CLIP ViT-B/32, and iSEARLE-XL, using CLIP ViT-L/14 as the backbone. From now on, we will refer to ViT-B/32 and ViT-L/14 as B/32 and L/14, respectively.
@@ V-A Implementation Details
[…]
Using a single NVIDIA A100 40GB GPU, iSEARLE-XL-OTI takes $\sim$35 seconds for a single image and $\sim$1.1 seconds per image with batch size 256. The training of $\phi$ for iSEARLE-XL takes 12 hours in total on a single A100 GPU. Throughout all the experiments, we adopt the pre-processing technique introduced in [5]. For retrieval, we normalize both the query and index set features to have a unit $L_{2}$-norm.
To generate the phrases used for the regularization with $\mathcal{L}_{gpt}$, we exploit the GPT-Neo-2.7B model with 2.7 billion parameters developed by EleutherAI. For each of the 20,932 class names of the Open Images V7 dataset [51], we generate 256 phrases a priori with a temperature of 0.5 and a maximum length constraint of 35 tokens. The whole process requires about 12 hours to execute on a single A100 GPU. Since this operation only needs to be performed once, the time requirements are manageable.
We use the same set of templates during both training and inference. Since the FashionIQ dataset provides two relative captions per triplet, at inference time, we concatenate them using the conjunction ``and". To ensure our method remains invariant to the concatenation order, we use both possible concatenation orders and average the resulting features.
@@ V-B Quantitative Results
We provide the results of both iSEARLE and iSEARLE-OTI. We compare our method with several zero-shot baselines: 1) Text-only: we compute the similarity using only the CLIP features of the relative caption; 2) Image-only: we retrieve the most similar images to the reference one; 3) Image + Text: we sum together the CLIP features of the reference image and the relative caption; 4) Captioning: we substitute the pseudo-word token with the caption of the reference image obtained via a pre-trained captioning model [62] [^4] https://huggingface.co/laion/CoCa-ViT-B-32-laion2B-s13B-b90k. In addition, we compare the proposed approach with the previous version of our method [9] and state-of-the-art ZS-CIR methods: Pic2Word [18], Context-I2W [33], and LinCIR [34]. For a fair comparison, we only consider competing methods that rely on the CLIP model without fine-tuning its weights and that, at inference time, do not require any additional pre-trained models, such as LLMs.
TABLE III: Quantitative results on CIRR test set. indicates results from the original paper. – denotes results not reported in the original paper.
[…]
@@
TABLE IV: Quantitative results on CIRCO test set. indicates results from the original paper.
[…]
@@
TABLE V: Quantitative results on CIRCO test set for each semantic category.
@@
CIRCO In Table IV, we report the results for the CIRCO test set. Firstly, we observe that, in contrast to FashionIQ and CIRR, Image+Text outperforms Image-only and Text-only. This result indicates that CIRCO contains queries where both the reference image and the relative caption are equally crucial for retrieving the target images. Secondly, iSEARLE achieves a considerable improvement over all the baselines and even outperforms Pic2Word, despite a smaller backbone. Considering the L/14 backbone, iSEARLE-XL would achieve the best results. However, we recall that the conference version of our method SEARLE-XL [9] was used to ease the annotation process of CIRCO. Therefore, since the core of the approach proposed in this work is the same, it is likely that the results could exhibit some sort of bias. Still, we report them for completeness.
@@
Domain Conversion Table VI illustrates the results for the domain conversion task. Following [18], the query images are sourced from 200 classes of ImageNet [58] validation set, while the target ones belong to ImageNet-R [60]. We recall that we relied on the ImageNet unlabeled test set as the pre-training dataset of our method, so there is no overlap with the evaluation images of the domain conversion task. The purpose of this experiment is to study how our model can convert the domain of a query image by using the prompt ``$\{domain\}$ of $S_{*}$", where $\{domain\}$ is a word that indicates the domain, e.g. toy or origami. We consider the retrieved image correct if its class is the same as that of the query image and its domain matches the one specified by the prompt. For instance, given the domain ``toy" and a query image containing a real-world shark, the goal is to retrieve images depicting a shark toy. The results show that, both for the B/32 and L/14 backbones, our method outperforms all the baselines.
TABLE VI: Quantitative results for the domain conversion task. The query images are from the ImageNet validation set, while the target ones belong to ImageNet-R. indicates results from the original paper.
@@
Object Composition We provide the results for the object composition task on the COCO validation set [17] in Tab. VII. This task aims to retrieve an image comprising an object specified with a single query image and other objects described with text. The object composition task closely resembles the personalized retrieval one [38], differing mainly in that the latter involves queries composed of multiple images depicting the same object instance. Following [18], we use the prompt ``a photo of $S_{*}$, $\{obj_{1}\}$ and $\{obj_{2}\}$, …, and $\{obj_{n}\}$", where $\{obj_{i}\}$ are text descriptions of objects, e.g. mouse, laptop or kite. For both the B/32 and L/14 backbones, we notice that our approach achieves state-of-the-art results. Considering the L/14 backbone, we observe a performance gap between the OTI and $\phi$ variants of the proposed method. We suppose this result is due to the similarity between the object composition and personalized retrieval tasks, where it has been shown that an optimization-based method is more suitable and achieves better performance [38]. The results obtained by PALAVRA [38] on the object composition task confirm our hypothesis, as it achieves significantly better relative performance than in composed image retrieval, e.g. on CIRCO. In addition, we observe that LinCIR [34] obtains considerably worse performance than iSEARLE-XL. We suppose this outcome is due to their language-only training strategy, which makes their model struggle to capture fine-grained visual details.
TABLE VII: Quantitative results for the object composition task on COCO validation set. indicates results from the original paper.
Backbone Method R@1@1@ 1 R@5@5@ 5 R@10@10@ 10
@@
Discussion Our experiments show that the proposed approach consistently achieves commendable performance across different datasets, highlighting its robustness. Compared to the baselines, our method has better generalization and adaptability capabilities to tasks beyond standard composed image retrieval, such as the domain conversion and object composition. Moreover, iSEARLE and iSEARLE-OTI obtain comparable results in most scenarios. This confirms the effectiveness of the distillation process, which offers a significant efficiency improvement without sacrificing the performance. Finally, the fine-grained evaluation on CIRCO reveals the intrinsic limitations of current ZS-CIR methods, which struggle when dealing with queries involving complex semantic aspects such as negation and viewpoint changes.